OpenStudy (anonymous):

reduce to lwest terms 1 2/3 + 9/5

OpenStudy (anonymous):

1 2/3 + 9/5 = 5 / 3 + 9 / 5 = 25 / 15 + 27 / 15 = 52 / 15